Darlas dad is 4 times older than Darla. Darlas mom is 3 times older that Darla.
The sum of all thier ages is 88. How do I find the age of each person??

Answer by stanbon(75887)

Darlas dad is 4 times older than Darla. Darlas mom is 3 times older that Darla.
The sum of all thier ages is 88.
-----
Equations:
f + d + m = 88
f = 4d
m = 3d
----
Substitute for "f" and "m" and solve for "d":
4d + d + 3d = 88
8d = 88
d = 11 (Darla's age now)
f = 4d = 44 (Father's age now)
m = 3d = 33 (Mom's age now)
===================
Cheers,
Stan H.
==============